The mass of a typical car is m=1000kg. so its weight on the Earth’s surface is mg=9800N. Suppose you have two containers, one with N electrons and another with N protons. These two containers are placed a distance 10m apart, and the electric force between them is equal to the weight of the car. Find N.1.      



1
Expert's answer
2021-04-26T17:12:01-0400

"N = \sqrt{\dfrac{9800\times100}{9\times10^9\times1.602\times{1.602\times10^{-38}}}}=6.154\times10^{16}"

Answer: 6.154 "\times" 1016.